# SABR-Informed Multitask Gaussian Process: A Synthetic-to-Real Framework for Implied Volatility Surface Construction

### v1

@@ Abstract
Constructing the Implied Volatility Surface (IVS) is a challenging task in quantitative finance due to the complexity of real markets and the sparsity of market data. Structural models like Stochastic Alpha Beta Rho (SABR) model offer interpretability and theoretical consistency but lack flexibility, while purely data-driven methods such as Gaussian Process regression can struggle with sparse data. We introduce SABR-Informed Multi-Task Gaussian Process (SABR-MTGP), treating IVS construction as a multi-task learning problem. Our method uses a dense synthetic dataset from a calibrated SABR model as a source task to inform the construction based on sparse market data (the target task). The MTGP framework captures task correlation and transfers structural information adaptively, improving predictions particularly in data-scarce regions. Experiments using Heston-generated ground truth data under various market conditions show that SABR-MTGP outperforms both standard Gaussian process regression and SABR across different maturities. Furthermore, an application to real SPX market data demonstrates the method’s practical applicability and its ability to produce stable and realistic surfaces. This confirms our method balances structural guidance from SABR with the flexibility needed for market data.
@@
Gaussian Process , Implied Volatility , Multi-Task Learning , SABR , Machine Learning.
@@ 1 Introduction
Constructing the Implied Volatility Surface (IVS) is an important and difficult task in quantitative finance. A common difficulty arises from sparse market option data, especially for long-dated expiration or strikes far from the money. This sparsity causes difficulties for conventional construction approaches.
@@
Structural models, such as the SABR model [1], are popular because they capture volatility smiles and skews with few parameters. The SABR model assume that the underlying asset and its volatility follow specific stochastic differential equations. While these models provide reasonable interpolations, their predetermined functional forms may suffer from a lack of flexibility and not fully reflect market complexities. In contrast, data-driven approaches offer more flexibility, adapting to market patterns without imposing rigid structures. Gaussian Processes (GP) [2], provide a Bayesian framework for regression. However, these methods typically require substantial data to achieve reliable results. With sparse observations, GP may fail to capture patterns of implied volatility.
@@
In order to address the challenges in data-driven approaches, recent works discuss how to incorporate machine learning methods with prior financial knowledge. Cousin et al. [3], Chataigner et al. [4], Ackerer et al. [5], Zheng et al. [6], Gonon et al. [7], Hoshisashi et al. [8] present methods that imposing constraints derived from financial theory within a neural network or a Gaussian Process. In contrast, Chen et al. [9] use typical transfer learning methods for option pricing. A neural network is trained on a structural model’s synthetic data (generated by the Black-Scholes model) is primarily, and it provide a good initialization for subsequent fine-tuning on empirical data.
@@
Different to their methods, we present the SABR-Informed Multi-Task Gaussian Process (SABR-MTGP) which takes advantage of both structural and data-driven models. The key idea is to treat the SABR-generated IVS as an information source (a ‘source’ task) and market observations as the primary target task. Then, a multi-task learning framework [10, 11] is used where the relationship between the theoretical structure (SABR) and empirical observations (market data) is learned adaptively during a joint optimization process. The implementation consists of two main stages. First, we generate a synthetic dataset using the calibrated SABR model. This dataset embodies smile and term-structure characteristics typical of SABR dynamics. Second, we train a multi-task Gaussian Process (MTGP) model simultaneously on these dense synthetic data and the sparse market observations. The MTGP framework learns correlation patterns between tasks through shared and task-specific covariance components. We use task embeddings with hierarchical regularization to appropriately balance how much structural guidance from SABR is used when constructing the representation from real market data. This avoids both overly rigid adherence to SABR patterns and complete disregard for its structural insights.
We evaluated our method through numerical experiments using the Heston model [12] to generate sparse ‘market’ datasets. Comparing SABR-MTGP against standard GP and SABR with interpolation, we find that our approach generally yields more accurate volatility surfaces across diverse market conditions and maturities, with particular advantages in sparse data regions. We further demonstrate the practical applicability of our method through a case study on real SPX market data. These results show the effectiveness of our method in integrating financial domain knowledge within a flexible machine learning framework.
@@
Gaussian processes have found increasing application across finance and economics due to their flexibility and ability to quantify uncertainty. A significant body of work focuses on derivative pricing, valuation, and hedging. GP models have been used for constructing implied volatility curves [13, 14], constructing financial term structures [15], accelerating pricing and hedging calculations for various options [13, 16, 17, 18, 19], modeling derivative portfolios for CVA computation [20], pricing complex insurance products like variable annuities [21], and approximating Greeks for hedging [22]. Beyond direct pricing, GP models serve as flexible tools for calibrating implied and local volatility surfaces [14, 4] or measuring portfolio tail risk [23].
Furthermore, GP models are employed in broader financial and economic time series analysis, including imputing missing financial data [24], forecasting real estate prices [25], modeling inflation dynamics [26], understanding determinants of carbon market prices [27], and developing nonparametric vector autoregressions for macroeconomic analysis [28].
While these studies highlight the versatility of GP models, our work specifically contributes by employing a multi-task learning framework informed by a structural model (SABR) to address the challenge of sparse data in IVS construction, learning the relationship between the structural prior and market data adaptively.
The remainder of the paper is organized as follows: Section 2 introduces some basic concept in implied volatility surface and Gaussian Process. Section 3 provides the details of the proposed methods. In Section 4, we describe the setup of the numerical experiments and the results are presented in Section 5. In Section 6, we demonstrate the performance of the proposed method in real market data.
2 Background
This section provides a necessary background on the implied volatility surface and the Gaussian process.
2.1 Implied Volatility Surface
The Black-Scholes model [29] is a basic framework for European option pricing under specific assumptions: key assumptons include that the underlying asset price follows the Geometric Brownian Motion, and asset volatility are constants. Observed market option prices frequently deviate from the prices given by the Black-Scholes model under the assumption of constant volatility. This difference is captured through the Implied Volatility (IV), a way to match the theoretical framework with market observations. For an observed market price of call options $C_{mkt}(K,\tau)$ with time-to-maturity $\tau$ and strike price $K$, the corresponding IV $\sigma_{mkt}(K,\tau)$ is the unique positive value that makes the following equation holds:
@@ 2.1 Implied Volatility Surface
where $C_{BS}$ is the Black-Scholes formula, and $\theta_{BS}$ is the collection of rest parameters in the Black-Scholes formula. Finding $\sigma_{mkt}(K,\tau)$ needs solving the Black-Scholes formula backwards, usually done with iterative root-finding algorithms (like Newton-Raphson). When $\sigma_{mkt}(K,\tau)$ is shown as a function of both variables, the resulting three-dimensional surface is the Implied Volatility Surface (IVS). The patterns of IVS reflect market participants’ risk preferences and expectations about future price dynamics beyond the standard Black-Scholes assumptions. Accurate IVS construction is important for derivative pricing, hedging strategies, and risk management across option portfolios.
2.2 Gaussian Process
A Gaussian Process (GP) is a Bayesian non-parametric approach that can be used in regression problems [2]. A GP defines a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that any finite set of function values has a joint Gaussian distribution. A GP is fully specified by its mean function $m:\mathbb{R}^{d}\to\mathbb{R}$ and covariance function (or kernel) $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$:
$f\sim\mathcal{GP}(m,k).$
The mean function is often set to zero for simplicity. The kernel encodes assumptions about function properties like smoothness or characteristic length-scales, defining the covariance between function values at any two input points. A commonly used kernel is the Matérn 5/2 kernel function which is a second-order differentiable kernel function:
[…]
$\text{for}\quad\mathbf{x,x^{\prime}}\in\mathbb{R}^{d}$, where $\sigma_{f}$ is the scaling hyperparameter and $\gamma$ is the length-scales hyperparameter.
Given training data $\mathcal{D}=\{(\mathbf{X,y})\}$ where the components in $\mathbf{y}=[y_{1},\dots,y_{N}]^{\intercal}$ are noisy observations at $N$ locations $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{N}]^{\intercal}$. $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ where $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{n}^{2})$ is noise. As a result, a predictive distribution for the function value $f_{\ast}$ at a new test point $\mathbf{x}_{\ast}$ is given by the posterior $p\big{(}f_{\ast}(\mathbf{x}_{\ast})|\mathcal{D}\big{)}\sim\mathcal{N}(\mu_{\ast},\sigma_{\ast})$, where
$\mu_{\ast}$ $=k(\mathbf{X,x_{\ast}})^{\intercal}\big{(}k(\mathbf{X,X})+\sigma_{n}^{2}\mathbf{I}\big{)}^{-1}\mathbf{y},$ (2)
$\sigma_{\ast}^{2}$ $=k(\mathbf{x}_{\ast},\mathbf{x}_{\ast})-k(\mathbf{X,x_{\ast}})^{\intercal}\big{(}k(\mathbf{X,X})+\sigma_{n}^{2}\mathbf{I}\big{)}^{-1}k(\mathbf{X,x_{\ast}}).$ (3)
Here, $k(\mathbf{X,x_{\ast}})\coloneqq\left[k(\mathbf{x}_{1},\mathbf{x_{\ast}}),\dots,k(\mathbf{x}_{N},\mathbf{x_{\ast}})\right]^{\intercal}$, $\mathbf{I}$ is an identity matrix and
$k(\mathbf{X,X})\coloneqq\begin{bmatrix}k(\mathbf{x}_{1},\mathbf{x}_{1})&\dots&k(\mathbf{x}_{1},\mathbf{x}_{N})\\ \vdots&\ddots&\vdots\\ k(\mathbf{x}_{N},\mathbf{x}_{1})&\dots&k(\mathbf{x}_{N},\mathbf{x}_{N})\end{bmatrix}.$
The kernel hyperparameters and noise variance $\sigma_{n}^{2}$ are usually estimated by maximizing the marginal likelihood of the data. This provides an automatic way to balance data fit and model complexity [30]. GP provide a Bayesian framework that can model complex, non-linear relationships without imposing rigid functional forms. However, they can face challenges when training data is sparse, as often happens with implied volatility observations.
3 SABR-Informed Multi-Task Gaussian Process
@@ 3.1 Problem Formulation
The primary objective of this work is to construct an accurate Implied Volatility Surface (IVS). The IVS, denoted as $\sigma(K,\tau)$, is a function of the option strike price $K$ and its time-to-maturity $\tau$. We can represent the input features as a two-dimensional vector $\mathbf{x}=[K,\tau]^{\intercal}$. The task is thus to learn a regression model $f:\mathbb{R}^{2}\to\mathbb{R}^{+}$ that maps these input features to the corresponding implied volatility, i.e., $y=f(\mathbf{x})=\sigma(K,\tau)$.
@@
To address this challenge, our methodology transfers information from a structural financial model, specifically the SABR model. The core idea is to generate a dense synthetic dataset, denoted $\mathcal{D}_{\mathcal{S}}$, using a calibrated SABR model. This synthetic dataset acts as a source of information to guide the learning process and prediction, especially in regions where market data is scarce. The subsequent sections will show how this synthetic data is generated and incorporated within a multi-task learning framework.
@@ 3.2 Synthetic Data Generation via SABR
The SABR model describes the dynamics of the forward price $F_{t}\coloneqq S_{t}e^{(r-q)(T-t)}$ (where $S_{0}$ is the current price of underlying asset, $r$ is the risk-free interest rate, $q$ is the dividend rate, $T$ is the maturity) and its instantaneous volatility $\alpha_{t}$ via stochastic differential equations:
@@
Let $\mathbb{T}_{\text{mkt}}=\{\tau_{(1)},\dots,\tau_{(P)}\}$ be the set of unique maturities in the market data $\mathcal{D}_{\mathcal{T}}$, ordered $\tau_{(1)}<\tau_{(2)}<\dots<\tau_{(P)}$. For each maturity slice $\tau_{(p)}$, we calibrated SABR parameters $(\alpha_{(p)},\rho_{(p)},\nu_{(p)})$ to match the observed market smile by minimizing the squared errors by Nelder-Mead algorithm [31]:
@@
Here, $F_{\tau_{(p)}}=S_{0}e^{(r-q)\tau_{(p)}}$ is the forward price at maturity $\tau_{(p)}$. We fixed $\beta=0.5$ throughout our experiments, a common choice for equity options. Calibrating all four SABR parameters ($\alpha,\beta,\rho,\nu$) simultaneously for each maturity slice can be challenging and numerically unstable, especially with sparse or noisy data. Calibrating $\beta$ often requires richer data across strikes than might be available, particularly for longer maturities. Fixing $\beta$ reduces the dimensionality of the optimization problem for each slice from four parameters to three ($\alpha,\rho,\nu$), significantly improving the stability and speed of the calibration process. The parameters $\alpha_{(p)}$, $\rho_{(p)}$, and $\nu_{(p)}$ were optimized within realistic ranges: $2.00.01<\alpha_{(p)}<$ 2.0, $-0.99<\rho_{(p)}<0.0$, and $1.50.05<\nu_{(p)}<$ 1.5. This yielded calibrated parameters $\{(\alpha_{(p)},\rho_{(p)},\nu_{(p)})\}_{p=1}^{P}$ for each observed maturity slice.
@@
  • If $\tilde{\tau}$ is between two calibrated maturities, $\tau_{(p)}<\tilde{\tau}<\tau_{(p+1)}$, we used piecewise linear interpolation to get parameters $(\alpha_{\tilde{\tau}},\rho_{\tilde{\tau}},\nu_{\tilde{\tau}})$. For example, for $\alpha$:
@@
3.3 Multi-Task Gaussian Process Formulation
To combine information from the dense synthetic data (SABR-generated) and sparse market observations, we used a Multi-Task Gaussian Process (MTGP) framework. Standard GP models model a single output, while MTGP models handle multiple related outputs by capturing correlations between tasks [10, 11].
@@ 3.3 Multi-Task Gaussian Process Formulation
For the shared components, we use learned task embeddings to infer the relationship. We modeled $h_{\mathcal{S}}$ and $h_{\mathcal{T}}$ as jointly Gaussian with a covariance structure:
@@
The term $\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}$ depends on task embeddings $\mathbf{e}_{\mathcal{Z}},\mathbf{e}_{\mathcal{Z}^{\prime}}\in\mathbb{R}^{d^{\prime}}$ ($d^{\prime}$ is a hyperparemeter that need to be predetermined) , a shared variance scaling parameter $\sigma^{2}_{h}$ and a length-scale parameter $l_{h}$. This allowed the model to learn the level of information sharing. Tasks with closer embeddings would have stronger correlation through the shared component. With components $g_{\mathcal{S}},g_{\mathcal{T}},h_{\mathcal{S}},h_{\mathcal{T}}$, the overall covariance structure follows an Intrinsic Coregionalization Model (ICM) [10]:
@@
Where $k(\mathbf{X}_{\mathcal{Z}},\mathbf{X}_{\mathcal{Z^{\prime}}})$ is the matrix of kernel evaluations between input points, and $\mathbf{1}_{R\times H}$ is an $R\times H$ matrix of ones.

### v2

@@ Abstract
This study introduces a SABR-informed multitask Gaussian process for constructing implied volatility surfaces from sparse option quotes. We treat a dense synthetic dataset generated by a calibrated SABR model as the source task and market option quotes as the target task. Within the multitask Gaussian process framework, we learn cross-task dependence via task embeddings with hierarchical regularization, enabling adaptive transfer of structural information. On Heston ground truth across ten market regimes and in a case study with SPX options, the model achieves lower error than the single-task Gaussian process and SABR at near-term maturities and remains competitive at long-term maturities, while satisfying standard no-arbitrage conditions. The framework combines the theory-driven structure with nonparametric Bayesian regression and yields reliable implied volatility surfaces for risk management.
@@
Gaussian Process , Implied Volatility , Multitask Learning , SABR , Machine Learning
journal: Journal of Computational Science
@@ 1 Introduction
Constructing the implied volatility surface (IVS) from sparse option quotes is a standard task in practice. Trading desks use the surface to support volatility quoting and the day-to-day risk management of vanilla option books across strikes and maturities. Structuring teams and model developers use it as an input to the calibration of downstream pricing models. The task is particularly challenging when market liquidity is limited, for example around the market open or for less liquid option products, and in illiquid regions such as far-from-the-money strikes and long maturities where observations are sparse and unevenly distributed. These operational requirements motivate surface construction methods that remain stable under data sparsity while producing smooth and no-arbitrage extrapolations in the wings.
@@
Structural models, such as the SABR model [23], are popular because they capture volatility smiles and skews with few parameters. These models assume that the underlying asset and its volatility follow specific stochastic differential equations. While these models provide reasonable interpolations, their predetermined functional forms may lack the flexibility to fully reflect market complexities. In contrast, data-driven approaches such as the Gaussian process (GP) [38], a nonparametric Bayesian method that can be used in IVS construction [15, 36], offer more flexibility. However, this adaptability requires sufficient data. With sparse observations, they can be prone to overfitting and may violate no-arbitrage conditions.
@@
In order to address the challenges in data-driven approaches, recent works discuss how to incorporate prior financial knowledge into machine learning methods. The literature [12, 7, 1, 41, 20, 27] presents methods that impose constraints derived from financial theory within a neural network or a Gaussian process. In contrast, Chen et al. [8] use transfer learning methods for option pricing. A neural network is trained on synthetic data generated from the Black-Scholes model, and it provides a good initialization for subsequent fine-tuning on empirical data.
@@
Inspired by the transfer learning approach of Chen et al. [8], we propose a SABR-informed multitask Gaussian process (SABR-MTGP). The key idea is to treat the SABR-generated IVS as an information source (a ‘source’ task) and market observations as the primary target task. Then, a multitask learning framework [5, 2] is used where the relationship between the theoretical structure (SABR) and empirical observations is learned adaptively during a joint optimization process. In our experiments, the proposed SABR-MTGP stabilizes data-sparse wings (far-from-the-money strikes) relative to a single-task GP, while correcting systematic misspecification that can arise from purely SABR models. The implementation consists of two main stages. First, we generate a synthetic dataset using the calibrated SABR model. This dataset embodies smile and term-structure patterns characteristic of SABR dynamics. Second, we train a multitask Gaussian process (MTGP) model simultaneously on these dense synthetic data and the sparse market observations. The MTGP framework learns the correlation between tasks through the shared covariance component. We use task embeddings with hierarchical regularization to appropriately balance how much structural guidance from SABR is used when constructing the IVS from real market data. SABR-MTGP improves IVS reconstruction from sparse quotes, but it is not a replacement for SABR-based dynamic hedging, or OTC derivatives valuation. Our scope is IVS construction from sparse observations. Extending to hedging and OTC derivatives valuation requires market-consistent dynamics (e.g., local or stochastic local volatility) and Greek validation. These are left to future work.
@@
Implied volatility surface construction can be organized into three complementary approaches. First, structural models specify dynamics that reproduce smiles and term structures with interpretable parameters, including SABR [23], Heston [25], Variance Gamma [32] and the double-exponential model [28]. More recently, rough volatility models [3, 18] emphasize fractional regularity in volatility paths. Second, parametric surface parameterizations impose shape and no-arbitrage structure, e.g., SVI [19], SSVI [17], and spline-based families [11]. Third, machine learning approaches, including the Gaussian process [15, 36, 7] and neural networks [1, 41, 20, 27, 34, 10] provide data-driven alternatives.
Relative to transfer learning approaches for options (Chen et al. [8]), our contribution is to implement the transfer learning in the Gaussian process via a task kernel with learnable embeddings and a hierarchical prior that adapts cross-task correlation.
Besides implied volatility surface construction, Gaussian processes have found increasing application across finance and economics due to their flexibility and ability to quantify uncertainty. A significant body of work focuses on derivative pricing, valuation, and hedging. GP models have been used for constructing financial term structures [13], accelerating pricing and hedging calculations for various options [15, 31, 21, 29, 26], modeling derivative portfolios for CVA computation [14], pricing complex insurance products like variable annuities [22], and approximating Greeks for hedging [30]. Beyond direct pricing, GP models serve as flexible tools for calibrating implied and local volatility surfaces [36, 7] or measuring portfolio tail risk [35]. Furthermore, GP models are employed in broader financial and economic time series analysis, including imputing missing financial data [16], forecasting real estate prices [39], modeling inflation dynamics [9], understanding determinants of carbon market prices [37], and developing nonparametric vector autoregressions for macroeconomic analysis [24].
The remainder of the paper is organized as follows: Section 2 introduces basic concepts of the implied volatility surface. Section 3 provides the details of the proposed methods. In Section 4, we describe the setup of the numerical experiments and the results are presented in Section 5. In Section 6, we demonstrate the application of the proposed methods using real market data.
2 Background: Implied Volatility Surface
The Black-Scholes model [4] is a basic framework for European option pricing under specific assumptions: the underlying asset price follows a geometric Brownian motion, and the asset’s volatility is constant. In practice, this constant volatility assumption is inconsistent with observed option prices, which typically exhibit patterns such as volatility smiles and skews. To reconcile the Black-Scholes formula with market prices, the concept of Implied Volatility (IV) is introduced. For an observed market price of call options $C_{mkt}(K,\tau)$ with time-to-maturity $\tau$ and strike price $K$, the corresponding IV $\sigma_{mkt}(K,\tau)$ is the unique positive value that makes the following equation hold:
@@ 2 Background: Implied Volatility Surface
where $C_{BS}$ is the Black-Scholes formula, and $\theta_{BS}$ is the collection of remaining parameters in the Black-Scholes formula. Computing $\sigma_{mkt}(K,\tau)$ requires inverting the Black-Scholes formula using iterative root-finding algorithms (such as Newton-Raphson). When $\sigma_{mkt}(K,\tau)$ is shown as a function of both variables, the resulting three-dimensional surface is the Implied Volatility Surface (IVS). The patterns of IVS reflect market participants’ risk preferences and expectations about future price dynamics beyond the standard Black-Scholes assumptions. Accurate IVS construction is important for derivative pricing, hedging strategies, and risk management across option portfolios.
3 SABR-Informed Multitask Gaussian Process
@@ 3.1 Problem Formulation
The objective of this work is to construct an accurate Implied Volatility Surface (IVS). The IVS is a function of the option strike price $K$ and its time-to-maturity $\tau$. We can represent the input features as a two-dimensional vector $\mathbf{x}=[K,\tau]^{\intercal}$. The task is to learn a regression model $f:\mathbb{R}^{2}\to\mathbb{R}^{+}$ that maps these input features to the corresponding implied volatility.
@@
To address this challenge, our methodology transfers information from a structural financial model, specifically the SABR model. The core idea is to generate a dense synthetic dataset, denoted $\mathcal{D}_{\mathcal{S}}$, using a calibrated SABR model. This synthetic dataset acts as a source of information to guide the learning process and prediction, especially in regions where market data is scarce. The subsequent sections will show how this synthetic data is generated and incorporated within a multitask learning framework.
@@ 3.2 Synthetic Data Generation via SABR
The SABR model describes the dynamics of the forward price $F_{t}\coloneqq S_{t}e^{(r-q)(T-t)}$ (where $S_{0}$ is the current price of the underlying asset, $r$ is the risk-free interest rate, $q$ is the dividend rate, $T$ is the maturity) and its instantaneous volatility $\alpha_{t}$ via stochastic differential equations:
@@
Let $\mathbb{T}_{\text{mkt}}=\{\tau_{(1)},\dots,\tau_{(P)}\}$ be the set of unique maturities in the market data $\mathcal{D}_{\mathcal{T}}$, ordered $\tau_{(1)}<\tau_{(2)}<\dots<\tau_{(P)}$. For each maturity slice $\tau_{(p)}$, we calibrated SABR parameters $(\alpha_{(p)},\rho_{(p)},\nu_{(p)})$ to match the observed market smile by minimizing the squared errors using the Nelder-Mead algorithm [33]:
@@
Here, $F_{\tau_{(p)}}=S_{0}e^{(r-q)\tau_{(p)}}$ is the forward price at maturity $\tau_{(p)}$. Calibrating all four SABR parameters ($\alpha,\beta,\rho,\nu$) simultaneously for each maturity slice can be challenging and numerically unstable, especially with sparse or noisy data. Calibrating $\beta$ often requires richer data across strikes than might be available, particularly for longer maturities. In practice, fixing the backbone parameter $\beta$ is not restrictive for smile fitting. The remaining SABR parameters $(\alpha,\rho,\nu)$ provide sufficient flexibility to reproduce the observed smile with high accuracy (Zhang and Fabozzi [40]), while improving calibration stability under sparse data. We fixed $\beta=0.5$ throughout our experiments. Robustness with respect to the choice of $\beta$ is assessed in Section 5.4 (Figure 8). The parameters $\alpha_{(p)}$, $\rho_{(p)}$, and $\nu_{(p)}$ were optimized within realistic ranges: $0.01<\alpha_{(p)}<2.0$, $0.99<\rho_{(p)}<0.0$, and $0.05<\nu_{(p)}<1.5$. This yielded calibrated parameters $\{(\alpha_{(p)},\rho_{(p)},\nu_{(p)})\}_{p=1}^{P}$ for each observed maturity slice.
@@
[…]
@@
3.3 Multitask Gaussian Process Formulation
To combine information from the dense synthetic data and sparse market observations, we used a Multitask Gaussian process (MTGP) framework.
A Gaussian process (GP) is a Bayesian nonparametric approach that can be used in regression problems [38]. A GP defines a distribution over functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that any finite set of function values has a joint Gaussian distribution. Standard GP models handle a single output, while MTGP models handle multiple related outputs by capturing correlations between tasks [5, 2].
@@ 3.3 Multitask Gaussian Process Formulation
For the shared components, we used learned task embeddings to infer the relationship. We modeled $h_{\mathcal{S}}$ and $h_{\mathcal{T}}$ as jointly Gaussian with a covariance structure:
@@
The term $\hat{C}_{\mathcal{Z},\mathcal{Z}^{\prime}}$ depends on task embeddings $\mathbf{e}_{\mathcal{Z}},\mathbf{e}_{\mathcal{Z}^{\prime}}\in\mathbb{R}^{d^{\prime}}$ ($d^{\prime}$ is a hyperparameter that needs to be predetermined), a shared variance scaling parameter $\sigma^{2}_{h}$ and a length-scale parameter $l_{h}$. This allowed the model to learn the level of information sharing. Tasks with closer embeddings would have stronger correlation through the shared component. With components $g_{\mathcal{S}},g_{\mathcal{T}},h_{\mathcal{S}},h_{\mathcal{T}}$, the overall covariance structure follows an Intrinsic Coregionalization Model (ICM) [5]:
@@
where $k(\mathbf{X}_{\mathcal{Z}},\mathbf{X}_{\mathcal{Z^{\prime}}})$ is the covariance matrix, and $\mathbf{1}_{R\times H}$ is an $R\times H$ matrix of ones.